I don't understand how to start this problem: David is Y years old. In 15 years, he will be two times as old as he was 10 years ago.
Answer by stanbon(75887)

David is Y years old.
In 15 years, he will be two times as old as he was 10 years ago.
Equation:
y+15 = 2(y-10)
y + 15 = 2y - 20
y = 35 yrs (David's age now)
------------
Cheers,
Stan H.
------------